# The Key Physics of Ice Premelting

### v1

@@ V Surface Intermolecular Forces
Here, $g_{sr}(h)$ is a short range contribution with exponentially fast decay which results from packing correlations,Chernov and Mikheev (1988); Henderson (1994); de Carvalho et al. (1994); Henderson (2005) short range attractive interactions,Derjaguin (1987); Israelachvili (1991) and finer effects such as renormalizable surface fluctuations.Chernov and Mikheev (1988); Henderson (2005) Additionally, $g_{vdw}(h)$ is a long range contribution that stems from the van der Waals interactions and decays in algebraic fashion.Parsegian and Ninham (1970); Parsegian (2005); Israelachvili (1991); Dietrich and Napiórkowski (1991)
@@ V.1 Packing and layering effects
Refer to caption Refer to caption
Figure 7: Calculation of disjoining pressures. a) Film thickness measured at solid-vapor equilibrium for the primary prism face as a function of temperature. b) The results are mapped to the corresponding Laplace pressure $\Delta p$ difference between bulk water and vapor at the same chemical potential to produce a disjoining pressure isotherm. Symbols are results from computer simulations for the TIP4P/Ice model. The lines are a fit to the asymptotic result of Eq.(10). Inset shows the calculated and predicted inverse susceptibilities as predicted from the second derivative of the interface potential, displaying regions of enhanced surface fluctuations. Reproduced with permission from Llombart et al.Llombart et al. (2020) Phys. Rev. Lett. 124 065702 (2020). Copyright ©2020 American Physical Society.
@@
Figure 7 shows an example for the disjoining pressure curve $\Pi=dg/dh$ as obtained from computer simulations using the TIP4P/Ice model of water.Abascal et al. (2005) The results are obtained by first measuring the equilibrium premelting thickness as a function of temperature (Fig.7-a), mapping this to the corresponding value of $\Delta p(T)$, as obtained accurately from thermodynamic integration, and plotting $\Delta p(T)$ as a function of $h_{e}(T)$ (Fig.7-b).Llombart et al. (2020)
The results show a monotonous decrease of the disjoining pressure. Therefore, the amplitude of the oscillations predicted in Eq.10 is too small to allow for the appearance of minima of the interface potential for films of less than several nanometers. In this fluctuation dominated regime, the oscillations predicted in the mean field approximation are washed away, and there is no longer a ’layering transition’ in the thermodynamic sense.Chernov and Mikheev (1988); Henderson (2005) However, the signature of underlying mean field transitions can survive as a flattening of the interface potential within given ranges of film thickness. In the computer simulations of Ref.Llombart et al. (2020), this rounded transitions could be observed as a marked oscillatory behavior of the inverse interfacial susceptibility (inset of Fig.7-b), which can be measured as the second derivative of the interface potential.
@@
To summarize, the emergence and gradual growth of premelting films, leading to a premelting film thickness of about 1 to 2 nm one Kelvin away from the triple point shown here is representative of all computer simulations studies.Conde, Vega, and Patrykiejew (2008); Limmer and Chandler (2014); Llombart et al. (2019, 2020); Llombart, Noya, and MacDowell (2020); Sibley et al. (2021); Baran et al. (2024) A detailed analysis suggests that in this range, there are enhanced surface fluctuations on growing from one bilayer to the next, but not a thermodynamic layering transition of the premelting film in the thermodynamic sense. Extrapolation of the fits beyond the range of film thicknesses achieved in simulation suggest that the oscillatory term of Eq.(10) could in principle lead to very shallow minima of the interface potential for the basal face,Llombart et al. (2020); Baran et al. (2024) but not for the prism face. However, the resulting minima are so shallow that they are very likely to vanish due to surface capillary waves. Under this assumption, the short range contribution of the interface potential is a monotonously decreasing function, which effectively promotes a situation of complete surface melting for both the basal and prism face, in a way similar to the prediction of the toy model of Eq.6.
@@ V.2 Van der Wals forces and the absence of complete melting
Unfortunately, it is not possible to assess directly the role of van der Waals forces by means of computer simulations for two reasons. The first and most important is that the van der Waals forces in computer simulations are cutoff beyond a truncation radius of about one nanometer, and this has been no exception in the study of ice premelting. The second is that, even if the interactions were calculated effectively, say, as in a lattice sum,Karasawa and Goddard (1989) the accuracy required to measure the effect of such forces would be likely prohibitive. Among other things, reaching a thick film beyond the nanometer requires approaching very carefully the triple point in the sub-Kelvin scale, but in practice, the uncertainty in the triple point of usual force fields is in the scale of one to two Kelvin.
@@
For most substances, the density of the solid phase is larger than that of the liquid phase, and the Hamaker constant is negative. In such case, the van der Waals contribution to the interface potential becomes positive, and favors complete surface melting. However, a well known anomaly of water is that the density of ice is smaller than that of liquid water. This means that the Hamaker constant ruling ice premelting is positive, a fine detail that was already noticed by Nozieres.Nozieres (1992) The implication is that the van der Waals forces take a negative value, and inhibit the formation of a thick wetting film. So, whatever is the behavior of the short range contribution, and whether it adopts minima due to the oscillations of the packing term or not, the contribution of the van der Waals forces is to inhibit the growth of the premelting film. As a result of this balance, the interface potential must have a minimum at intermediate distances, leading to a case of incomplete surface melting.
@@ V.4 Disgression: Into the realm of quantum electrodynamics
Figure 9: Hamaker functions as predicted by DLP theory. Blue lines are results for $A(h)$ as calculated from an accurate parametrization of the dielectric response of water and ice by Luengo et al.Luengo-Márquez, Izquierdo-Ruiz, and MacDowell (2022). For small $h$, $A(h)$ is positive as predicted by Eq.(12), and inhibits surface melting. For large $h$, $A(h)$ changes sign, becomes negative and favor surface melting. The inset shows the sign reversal occurs at about 760 nm. Red lines are results from Elbaum and Schick, which suffer from a poor parametrization of the dielectric response.Elbaum and Schick (1991) The sign of $A(h)$ is correct for films of thickness larger than 3 nm. The red dashed lines are results from Limmer and Chandler.Limmer and Chandler (2014) Adapted from Luengo et al. J. Chem. Phys.Luengo-Márquez, Izquierdo-Ruiz, and MacDowell (2022) 157 044704 (2022), with the permission of AIP Publishing.
The minimal model discussed above describes the qualitative behavior of ice premelting for film thicknesses less than hundreds of nanometers. Beyond this range, intermolecular forces become dominated by very subtle quantum electrodynamic effects, that are usually ignored altogether.Israelachvili (1991) Indeed, as emphasized in Eq.(11), the Hamaker ’constant’ is actually not a constant, but displays a very rich and complex behavior as the thickness of the premelting film becomes large. The full quantitative behavior of $A(h)$ can only be determined within the framework of Dzyaloshinskii-Lifshitz-Pitaevskii theory (DLP),Dzyaloshinskii, Lifshitz, and Pitaevskii (1961) which has been used extensively by Bostrom and collaborators to address a number of related problems in ice premelting.Boström et al. (2017, 2019); Esteso et al. (2020); Luengo-Márquez and MacDowell (2021); Li et al. (2022); Boström et al. (2023) Unfortunately, DLP theory is not amenable to an easy qualitative interpretation, and requires involved numerical calculations.Parsegian (2005) For this reason, a simple analytical approximation was worked out recently solely for this purpose.MacDowell (2019)
@@
where $\hbar$ is Planck’s constant, $c$ is the speed of light, $\nu_{\infty}\approx 0.01$ nm-1 is a wave-number corresponding to the principal electronic transition of water, $n_{i}$ are refractive indexes in the visible, and finally, $f(h)$ is a crossover function which remains close to unity for all $h\nu_{\infty}\ll 1$. In this regime of small $h$, the above equation accounts for correlated electronic dipole fluctuations as dictated by the material’s refractive indexes, in analogy with Eq.(12). As an anomaly of water, the refractive index of the solid phase is smaller than that of the liquid phase, $n_{s}<n_{l}$, which results in a positive Hamaker constant, exactly as predicted by Eq.(12).
@@
Figure 9 shows the full $h$ dependence of the Hamaker function with a recent reliable parametrization of the dielectric properties,Luengo-Márquez, Izquierdo-Ruiz, and MacDowell (2022) and compares it to previous calculations.Elbaum and Schick (1991); Limmer and Chandler (2014) The results of Luengo et al. are in qualitative agreement with Fiedler et al.Fiedler et al. (2020) (not shown), but Bostrom et al. have ever since opted for the parametrization of dielectric properties presented in Ref.Luengo-Márquez, Izquierdo-Ruiz, and MacDowell (2022).
@@ VI Water droplets on ice
Refer to caption Refer to caption Refer to caption
@@
The presence of water droplets has ever since been reported by a number of studies through the years,Ketcham and Hobbs (1969); Elbaum (1991); Elbaum, Lipson, and Dash (1993) including photographs of water droplets on the basal surface of hexagonal ice micro-crystals,Gonda, Arai, and Sei (1999) and large millimeter drops on polycrystalline ice.Makkonen (1997); Thiévenaz, Josserand, and Séon (2020); Demmenie et al. (2025); Sarlin et al. (2025) However, it was not until ten years that the quality of the reported pictures has allowed to appreciate this finding widely.Sazaki et al. (2012); Asakawa et al. (2016); Murata et al. (2016); Sazaki et al. (2022)
@@
Paradoxically, the experimental observation of liquid droplets on the ice surface has often been interpreted as illustrating the complete lack of equilibrium ice premelting,Knight (1996); Makkonen (1997); Sazaki et al. (2012); Asakawa et al. (2016); Murata et al. (2016); Sazaki et al. (2022); Demmenie et al. (2025) which is obviously at odds with experimental measurements of film thickness.Elbaum, Lipson, and Dash (1993); Dosch, Lied, and Bilgram (1995); Sadtchenko and Ewing (2002); Mitsui and Aoki (2019) This conflict emerges from a misunderstanding of Young’s equation, where it is often assumed that $\gamma_{{sv}}$ accounts for the surface tension of a bare interface between the unrelaxed ice surface and its vapor. At high temperature, the equilibrium solid-vapor interface can exhibit a significant amount of surface disorder consistent with the existence of a premelting film, and $\gamma_{{sv}}$ may become significantly different from the expected value for a perfect unrelaxed interface obtained by cleavege of the solid phase.
@@
The confusion is likely to result from an alternative expression for the surface free energy which is often found in the literature:Dash, Rempel, and Wettlaufer (2006); Nagata et al. (2019)
@@
In this expression, one can easily relate $I(h)$ with the interface potential, $g(h)$, whereupon, both expressions are fully equivalent. However, assuming that $I(h)=0$ for $h=0$ from the outset,Dash, Rempel, and Wettlaufer (2006); Nagata et al. (2019) leads to the misleading interpretation that the solid-vapor interface need be understood as a bare ice-vapor interface with no premelting whatsoever. Actually, the above equation is completely equivalent to Eq.(7), with the understanding that $I(h)=0$ holds not for $h=0$, but for $h=h_{e}$.
@@ VII Is the observation of surface steps consistent with ice premelting?
Figure 11: Experimental observation of the ice basal face in the neighborhood of the triple point. a) At slight supersaturation, elementary steps appear and spread horizontally. b) At higher saturation, droplets condense on the surface. c) At yet higher saturation, thick premelting films can appear below the droplets. d) Sketch showing, from left to right, the sequence of events observed in a), b) and c), as interpreted in Ref.. e) The ice phase diagram, illustrating the phase boundary lines in black, and the metastable prolongation of the condensation line in dashed black. The red line shows the region where condensation of water droplets first appears. The blue line shows the region where a second wetting film below the liquid droplets first appears. Notice the similarity with the sketch of Figure 4 predicted from the theory described in Section IX. Figures a), b), c) and d) reproduced with permission from Sazaki et al.Sazaki et al. (2012) PNAS 109 1052 (2012). Figure e) reproduced with permission from Murata et al.Murata et al. (2016) PNAS 113 E6741 (2016).
@@
Indeed, advances in optical microscopy and imaging have allowed in the last decade to observe (and report) the ice surface with unprecedented detail.Sazaki et al. (2013); Murata et al. (2016); Sazaki et al. (2022) Particularly, experiments performed by the Hokaido team have clearly revealed that the basal ice surface at exceedingly small supersaturation exhibits elementary steps in perfect agreement with the expected lattice structure, and results in crystal growth by terrace spreading or spiral growth.Sazaki et al. (2013); Murata et al. (2016); Sazaki et al. (2022) But the general wisdom in crystal growth theory is that this kind of smooth surfaces are in a highly ordered, low temperature state.Safran (1994); Chaikin and Lubensky (1995); Feigelson (2004); Akutsu and Yamamoto (2015) The question then arises, how can one reconcile the step formation expected for a highly ordered surface, with the highly disordered state expected for a premelted surface? Sazaki and collaborators interpreted their own observation as evidence for the lack of premelting layers on the ice surface, most noticeably, in Ref.Asakawa et al. (2016), but have thereafter sometimes stand by,Sazaki et al. (2022) sometimes moderated this claim.Murata, Nagashima, and Sazaki (2019); Nagata et al. (2019)
However, a recent combination of simple models borrowed from crystal growth theory and capillary wave theory, shows that the observation of significant premelting and step formation are both mutually acceptable and can be reconciled.Benet et al. (2016, 2019); Sibley et al. (2021)

### v2

@@ V Surface Intermolecular Forces
Here, $g_{sr}(h)$ is a short range contribution with exponentially fast decay which results from packing correlations,Chernov and Mikheev (1988); Henderson (1994); de Carvalho et al. (1994); Henderson (2005) short range attractive interactions,Deryagin (1987); Israelachvili (1991) and finer effects such as renormalizable surface fluctuations.Chernov and Mikheev (1988); Henderson (2005) Additionally, $g_{vdw}(h)$ is a long range contribution that stems from the van der Waals interactions and decays in algebraic fashion.Parsegian and Ninham (1970); Parsegian (2005); Israelachvili (1991); Dietrich and Napiórkowski (1991)
@@ V.1 Packing and layering effects
[…]
Figure 7: Calculation of disjoining pressures. a) Film thickness measured at solid-vapor equilibrium for the primary prism face as a function of temperature. b) The results are mapped to the corresponding Laplace pressure difference, $\Delta p$, between bulk water and vapor at the same chemical potential to produce a disjoining pressure isotherm. Symbols are results from computer simulations for the TIP4P/Ice model. The lines are a fit to the asymptotic result of Eq.(10). The inset shows inverse susceptibilities as calculated from the simulations (symbols) and predicted from the second derivative of the interface potential (lines), displaying regions of enhanced surface fluctuations. Reproduced with permission from Llombart et al.Llombart et al. (2020) Phys. Rev. Lett. 124 065702 (2020). Copyright ©2020 American Physical Society.
@@
Figure 7 shows an example for the disjoining pressure curve $\Pi=dg/dh$ of the prism face as obtained from computer simulations using the TIP4P/Ice model of water.Abascal et al. (2005) The curve is computed by first measuring the equilibrium premelting thickness as a function of temperature (Fig.7-a), mapping this to the corresponding value of $\Delta p(T)$, as obtained accurately from thermodynamic integration, and plotting $\Delta p(T)$ as a function of $h_{e}(T)$ (Fig.7-b).Llombart et al. (2020)
The results show a monotonous decrease of the disjoining pressure. Therefore, the amplitude of the oscillations predicted in Eq.10 is too small to allow for the appearance of minima of the interface potential for films of less than several nanometers. In this fluctuation dominated regime, the oscillations predicted in the mean field approximation are washed away, and there is no longer a ‘layering transition’ in the thermodynamic sense.Chernov and Mikheev (1988); Henderson (2005) However, the signature of underlying mean field transitions can survive as a flattening of the interface potential within given ranges of film thickness. In the computer simulations of Ref.Llombart et al. (2020), these rounded transitions could be observed as a marked oscillatory behavior of the inverse interfacial susceptibility (inset of Fig.7-b), which can be measured as the second derivative of the interface potential.
@@
To summarize, all computer simulation studies support the emergence and gradual growth of premelting films, with a film thickness of about 1 to 2 nm one Kelvin away from the triple point. Conde, Vega, and Patrykiejew (2008); Limmer and Chandler (2014); Llombart et al. (2019, 2020); Llombart, Noya, and MacDowell (2020); Sibley et al. (2021); Baran et al. (2024) A detailed analysis suggests that in this range, there are enhanced surface fluctuations on growing from one bilayer to the next, but not a thermodynamic layering transition of the premelting film in the thermodynamic sense. Extrapolation of the fits beyond the range of film thicknesses achieved in simulation suggest that the oscillatory term of Eq.(10) could in principle lead to very shallow minima of the interface potential for the basal face,Llombart et al. (2020); Baran et al. (2024) but not for the prism face. However, the resulting minima are so shallow that they are very likely to vanish due to surface capillary waves. Under this assumption, the short range contribution of the interface potential is a monotonously decreasing function, which effectively promotes a situation of complete surface melting for both the basal and prism face, in a way similar to the prediction of the toy model of Eq.6.
@@ V.2 Van der Wals forces and the absence of complete melting
Unfortunately, it is not possible to assess directly the role of van der Waals forces by means of computer simulations for two reasons. The first and most important is that the van der Waals forces in computer simulations are cutoff beyond a truncation radius of about one nanometer, and this has been no exception in the study of ice premelting. The second is that, even if the interactions were calculated effectively, say, as in a lattice sum,Karasawa and III (1989) the accuracy required to measure the effect of such forces would be likely prohibitive. Among other things, reaching a thick film beyond the nanometer requires approaching very carefully the triple point in the sub-Kelvin scale, but in practice, the uncertainty in the triple point of usual force fields is in the scale of one to two Kelvin.
@@
For most substances, the density of the solid phase is larger than that of the liquid phase, and the Hamaker constant is negative. In such case, the van der Waals contribution to the interface potential becomes positive, and favors complete surface melting. However, a well known anomaly of water is that the density of ice is smaller than that of liquid water. This means that the Hamaker constant ruling ice premelting is positive, a fine detail that was already noticed by Nozieres.Nozieres (1992) The implication is that the van der Waals forces take a negative value, and inhibit the formation of a thick wetting film. So, whatever is the behavior of the short range contribution, and whether it adopts minima due to the oscillations of the packing term or not, the contribution of the van der Waals forces is to inhibit the growth of the premelting film. As a result of this balance, the interface potential must have a minimum at intermediate distances, leading to a case of incomplete surface melting. Further notice that this behavior is generic and arises from the nature of the bulk ice and water phases, so that, to leading order, it does not depend on the nature of the exposed surface plane .
@@ V.4 Disgression: Into the realm of quantum electrodynamics
Figure 9: Hamaker functions as predicted by DLP theory. Blue lines are results for $A(h)$ as calculated from an accurate parametrization of the dielectric response of water and ice by Luengo et al.Luengo-Márquez, Izquierdo-Ruiz, and MacDowell (2022). For small $h$, $A(h)$ is positive as predicted by Eq.(12), and inhibits surface melting. For large $h$, $A(h)$ changes sign, becomes negative and favor surface melting. The inset shows the sign reversal occurs at about 760 nm. Orange lines are results from Elbaum and Schick, which suffer from a poor parametrization of the dielectric response.Elbaum and Schick (1991) The sign of $A(h)$ is correct for films of thickness larger than 3 nm. The red dashed lines are results from Limmer and Chandler.Limmer and Chandler (2014) Adapted from Luengo et al. J. Chem. Phys.Luengo-Márquez, Izquierdo-Ruiz, and MacDowell (2022) 157 044704 (2022), with the permission of AIP Publishing.
The minimal model discussed above describes the qualitative behavior of ice premelting for film thicknesses less than hundreds of nanometers. Beyond this range, intermolecular forces become dominated by very subtle quantum electrodynamic effects, that are usually ignored altogether.Israelachvili (1991) Indeed, as emphasized in Eq.(11), the Hamaker ’constant’ is actually not a constant, but displays a very rich and complex behavior as the thickness of the premelting film becomes large. The full quantitative behavior of $A(h)$ can only be determined within the framework of Dzyaloshinskii-Lifshitz-Pitaevskii theory (DLP),Dzyaloshinskii, Lifshitz, and Pitaevskii (1961) which has been used extensively by Bostr ö m and collaborators to address a number of related problems in ice premelting.Boström et al. (2017, 2019); Esteso et al. (2020); Luengo-Márquez and MacDowell (2021); Li et al. (2022); Boström et al. (2023) Unfortunately, DLP theory is not amenable to an easy qualitative interpretation, and requires involved numerical calculations.Parsegian (2005) For this reason, a simple analytical approximation was worked out recently solely for this purpose.MacDowell (2019)
@@
where $\hbar$ is Planck’s constant, $c$ is the speed of light, $\nu_{\infty}\approx 0.01$ nm-1 is a wave-number corresponding to the principal electronic transition of water, $n_{i}$ are refractive indexes in the visible, and finally, $f(h)$ is a crossover function which remains close to unity for all $h\nu_{\infty}\ll 1$. In this regime of small $h$, the above equation accounts for correlated electronic dipole fluctuations as dictated by the material’s refractive indexes. As an anomaly of water, the refractive index of the solid phase is smaller than that of the liquid phase, $n_{s}<n_{l}$, which results in a positive Hamaker constant, exactly as predicted by Eq.(12).
@@
Figure 9 shows the full $h$ dependence of the Hamaker function with a recent reliable parametrization of the dielectric properties,Luengo-Márquez, Izquierdo-Ruiz, and MacDowell (2022) and compares it to previous calculations.Elbaum and Schick (1991); Limmer and Chandler (2014) The results of Luengo et al. are in qualitative agreement with Fiedler et al.Fiedler et al. (2020) (not shown), but Bostr ö m et al. have ever since opted for the parametrization of dielectric properties presented in Ref.Luengo-Márquez, Izquierdo-Ruiz, and MacDowell (2022).
@@ VI Water droplets on ice
[…]
@@
The presence of water droplets has ever since been confirmed by a number of studies through the years,Ketcham and Hobbs (1969); Elbaum (1991); Elbaum, Lipson, and Dash (1993) including photographs of water droplets on the basal surface of hexagonal ice micro-crystals,Gonda, Arai, and Sei (1999) and large millimeter drops on polycrystalline ice.Makkonen (1997); Thiévenaz, Josserand, and Séon (2020); Demmenie et al. (2025); Sarlin et al. (2025) However, it was not until the last ten years that the quality of the reported pictures has allowed to appreciate this finding widely.Sazaki et al. (2012); Asakawa et al. (2016); Murata et al. (2016); Sazaki et al. (2022)
@@
Paradoxically, the experimental observation of liquid droplets on the ice surface has often been interpreted as illustrating the complete lack of equilibrium ice premelting,Knight (1996); Makkonen (1997); Sazaki et al. (2012); Asakawa et al. (2016); Murata et al. (2016); Sazaki et al. (2022); Demmenie et al. (2025) which is obviously at odds with experimental measurements of film thickness.Elbaum, Lipson, and Dash (1993); Dosch, Lied, and Bilgram (1995); Sadtchenko and Ewing (2002); Mitsui and Aoki (2019) This conflict emerges from a misunderstanding of Young’s equation, where it is often assumed that $\gamma_{{sv}}$ accounts for the surface tension of a bare interface between the unrelaxed ice surface and its vapor. At high temperature, the equilibrium solid-vapor interface can exhibit a significant amount of surface disorder consistent with the existence of a premelting film, and $\gamma_{{sv}}$ may become significantly different from the expected value for a perfect unrelaxed interface obtained by cleavage of the solid phase.
@@
The confusion is likely to result from an alternative convention for the surface free energy which is often found in the literature:Dash, Rempel, and Wettlaufer (2006); Nagata et al. (2019)
@@
Comparing this result with Eq.(7) , one can easily relate $I(h)$ with the interface potential, $g(h)$, whereupon, both results are fully equivalent. However, assuming that $I(h)=0$ for $h=0$ from the outset,Dash, Rempel, and Wettlaufer (2006); Nagata et al. (2019) leads to the misleading interpretation that the solid-vapor interface need be understood as a bare ice-vapor interface with no premelting whatsoever. Actually, the above equation is completely equivalent to Eq.(7), with the understanding that $I(h)=0$ holds not for $h=0$, but for $h=h_{e}$.
@@ VII Is the observation of surface steps consistent with ice premelting?
Figure 11: Experimental observation of the ice basal face in the neighborhood of the triple point. A) At slight supersaturation, elementary steps appear and spread horizontally. B) At higher saturation, droplets condense on the surface. C) At yet higher saturation, thick premelting films can appear below the droplets. D) Sketch showing, from left to right, the sequence of events observed in A), B) and C), as interpreted in Ref.Sazaki et al., 2012. E) The ice phase diagram, illustrating the phase boundary lines in black, and the metastable prolongation of the condensation line in dashed black. The red line shows the region where condensation of water droplets first appears. The blue line shows the region where a second wetting film below the liquid droplets first appears. Notice the similarity with the sketch of Figure 4 predicted from the theory described in Section IX. Figures A, B, C and D reproduced with permission from Sazaki et al.Sazaki et al. (2012) PNAS 109 1052 (2012). Figure E reproduced with permission from Murata et al.Murata et al. (2016) PNAS 113 E6741 (2016).
@@
In experiments performed by the Hokaido team, atomically smooth ice surfaces were prepared by epitaxial deposition from pure water vapor. This careful sample preparation, together with advanced optical microscopy have allowed in the last decade to observe (and report) the ice surface with unprecedented detail.Sazaki et al. (2013); Murata et al. (2016); Sazaki et al. (2022) Particularly, it was shown that the basal ice surface at exceedingly small supersaturation exhibits elementary steps in perfect agreement with the expected lattice structure, and results in crystal growth by terrace spreading or spiral growth.Sazaki et al. (2013); Murata et al. (2016); Sazaki et al. (2022) But the general wisdom in crystal growth theory is that this kind of smooth surfaces are in a highly ordered, low temperature state.Safran (1994); Chaikin and Lubensky (1995); Feigelson (2004); Akutsu and Yamamoto (2015) The question then arises, how can one reconcile the step formation expected for a highly ordered surface, with the highly disordered state expected for a premelted surface? Sazaki and collaborators interpreted their own observation as evidence for the lack of premelting layers on the ice surface, most noticeably, in Ref.Asakawa et al. (2016), but have thereafter sometimes stand by,Sazaki et al. (2022) sometimes moderated this claim.Murata, Nagashima, and Sazaki (2019); Nagata et al. (2019)
However, a recent combination of simple models borrowed from crystal growth theory and capillary wave theory, shows that the observation of significant premelting and step formation are both mutually acceptable and reconcilable.Benet et al. (2016, 2019); Sibley et al. (2021)